Foto Company makes 40,000 units per year of a part it uses in the products
it manufactures. The cost per unit of this part is shown below:

direct materials .............. $12.00
direct labor ..................  10.10
variable overhead .............   5.60
allocated fixed overhead ......  10.50
total ......................... $38.20

An outside supplier has offered to sell Foto Company 40,000 of these parts
for $36.70 per unit. If the company accepts this offer, the facilities now
being used to make the part could be used to make more units of a product
that is in high demand. The additional contribution margin earned on this
other product would be $310,000 per year.

Calculate the selling price per unit charged by the outside supplier that
would make Foto economically indifferent between making and buying the part.
Answer #1

Indifferent point

Manufacturing cost = Purchase cost

27.70*40000+310000 = 40000X

Selling price per unit charged by outside supplier = (1108000+310000)/40000 = 35.45 per unit